# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@
Abstract.
We present a self-contained treatment of the fundamental role that quasi-convex functions play in general (nonlinear second order) potential theories, which concerns the study of generalized subharmonics associated to a suitable closed subset (subequations) of the space of $2$-jets. Quasi-convex functions build a bridge between classical and viscosity notions of solutions of the natural Dirichlet problem in any potential theory. Moreover, following a program initiated by Harvey and Lawson in [23], a potential-theoretic approach is widely being applied for treating nonlinear partial differential equations (PDEs). This viewpoint revisits the conventional viscosity approach to nonlinear PDEs [13] under a more geometric prospective inspired by Krylov [36] and takes much insight from classical pluripotential theory. The possibility of a symbiotic and productive relationship between general potential theories and nonlinear PDEs relies heavily on the class of quasi-convex functions, which are themselves the generalized subharmonics of a pure second order constant coefficient potential theory.
@@ Introduction
This work is dedicated to foundational aspects of general (nonlinear) second order potential theories as initiated by Reese Harvey and Blaine Lawson, beginning with a trio of papers, [21], [22] and [23], published in 2009. A general potential theory concerns the study of generalized subharmonic functions with respect to a given constraint set on the space of $2$-jets, called a subequation. The boundary of a subequation provides a second order partial differential equation. Our aim is to give a comprehensive presentation of the main tools in quasi-convex analysis which are useful in order to study general potential theories as well as to deal with nonlinear PDEs by a the potential-theoretic approach of Harvey and Lawson, which (in many cases) can recast the operator theory in a flexible and more geometric way through the use of an associated potential theory. Recently, a powerful correspondence principle has been established in the presence of sufficient monotonicity for the constant coefficient case in the monograph [8] and in the additional presence of fiberegularity for the general case in the paper [11]. This correspondence principle is used to embed a given partial differential operator into an associated potential theory and then to deduce properties about the operator theory from the corresponding potential theoretic properties. Moreover, these properties are shared by any operator compatible to the given potential theory. In that light, the present work can be viewed as a prequel to the aforementioned works and our main aim is to provide a comprehensive introduction to the foundations on which the Harvey-Lawson theory rests. Given the comprehensive nature of the work, some of the results we present will be well known. Many others are much less well known, and this work aims to change that fact. There will also be new results and new interpretations of recent known results, which derive from the insight of Harvey and Lawson.
Before continuing with the presentation, an important remark concerning terminology should be made. We follow the usage of Harvey and Lawson (as well as [8] and [11]) by saying that an upper semicontinuous function $u$ is quasi-convex if a quadratic perturbation $u+\frac{\lambda}{2}|\cdot|^{2}$ is convex for some $\lambda\geqslant 0$, which is equivalent to asking that the Hessian of $u$ is bounded below (in the viscosity sense). This is often called semiconvexity (especially in the conventional viscosity literature such as [13]). We prefer the term quasi-convex since these functions are not “half-convex” in the same way that semicontinuous functions are “half-continuous” and the term quasi-convex is consistent with the use of quasi-plurisubharmonic in several complex analysis, as noted in [32]. Moreover, one should also note that in convex analysis, quasi-convexity is often used to mean that inverse images of half-lines bounded above are convex sets. Perhaps a different term such as “convexable” might be advisable, or to use Hessian bounded below as in [43]. Indeed, any $C^{2}$ function is locally quasi-convex, which shows that quasi-convex functions are far from being “half-convex”.
In order to set the stage for the role that quasi-convex functions play in general potential theories, we begin with a description of the program initiated by Harvey and Lawson. Their program was in part inspired by the geometric approach of Krylov [36] in order to formulate a general notion of ellipticity. In the pure second order case, Krylov calls a differential operator $u\mapsto F(D^{2}u)$ elliptic if it there exists an elliptic set for $F$; that is, an open subset $\Theta\subset\mathcal{S}(n)$ such that $F$ is increasing for $A\in\Theta$ and the differential inclusion $D^{2}u(x)\in\partial\Theta$ for each $x\in\Omega$ is called an elliptic branch of the equation $F(D^{2}u)=0$ if $\partial\Theta\subset\{F=0\}$. One can then “forget” about the particular form of the operator $F$ and concentrate instead on the constraint set $\Theta$ as many operators $F$ will give rise to the same $\Theta$. Harvey and Lawson recognized similarities of this idea in pluripotential theory and starting instead from closed sets $\mathcal{F}\subset\mathcal{S}(n)$ (the closure of $\Theta$), formulated a suitable notion of duality which recasts the notion of viscosity solutions in a precise topological framework and developed a monotonicity–duality method for proving the comparison principle in general potential theories. With additional work, this yields the comparisn principle for any fully nonlinear second-order partial differential which is compatible with given potential theory. The comparison principle of course implies uniqueness of solutions for the associated Dirichlet problem in both contexts (potential theory and PDEs). Another major breakthrough in the Harvey-Lawson approach is to use asymptotic interiors of the constraint set to determine the correct notion of boundary pseudoconvexity for which one can construct the barriers needed to complete Perron’s method for proving existence of solutions for the Dirichlet problem.
After their pioneering paper [23] for pure second order constant coefficient potential theories, a sequence of ambitious papers of Harvey and Lawson followed, such as [24] which studies the Dirichlet problem for fully nonlinear second-order equations on Riemannian manifolds, [27] which characterizes the validity of the strong maximum principle, [28] and [31] which treat the special Lagrangian potential equation, [21] on potential theory in calibrated geometry and [29] on the inhomogeneous Dirichlet problem for “natural” operators. Inspired by [23], Marco Cirant and the first author began an investigation on embedding large classes of differential operators into the potential theoretic framework of Harvey and Lawson. The variable coefficient pure second order and gradient-free cases were treated in [9] and [10] respectively, where the continuity of the fiber map $x\mapsto\mathcal{F}_{x}=\Theta(x)$ (now called fiberegularity) was shown to be a sufficient additional condition in order to transport the monotonicity-duality method from the constant coefficient to the variable coefficient setting. In addition, a rudimentary version of the correspondence principle was established in [9] and [10] but followed more explicitly the ideas of Krylov ((proper) elliptic maps $\Theta$ and associated branches of the PDE) as opposed to the subequation formulation. Joining forces in [8], a complete and robust treatment of constant coefficient potential theories and PDEs in Euclidian spaces by the monotonicity-duality method was given, including the aforementioned correspondence principle, a classification of monotonicity cone subequations and numerous illustrations of the method. Finally, the general variable coefficient case in the presence of sufficient monotonicity (proper ellipticity and directionality) and fiberegularity is treated by the authors and Cirant in [11].
Having recalled the context of the prequel that will be presented here, some declarations of intent are in order. The present work was born with the purpose of unifying and harmonizing the two articles [25] and [26] and of Harvey and Lawson on quasi-convex functions and their use in general potential theories, which were deposited in the arXiv in 2016. These two preprints are full of foundational aspects of non-smooth analysis, some well known, others less well known and many new things. In particular, the two unpublished manuscripts contain many refinements and efficient use of quasi-convex functions for unifying and simplifying many aspects of the conventional viscosity theory, and as such, merit a comprehensive and unified presentation. The authors of this work are extremely grateful to Harvey and Lawson for their encouragement to attempt this synthesis, which was the basis of the Master’s Thesis of the second author, supervised by the first author.
It should be noted that the present work is more than a mere merging of the two preprints [25] and [26], whose combined length is less than one third of the present work. Beyond the question of length, the current presentation benefits from the numerous developments in the theory over the past decade. The last three paragraphs of the introduction will indicate what is “new” here. Finally, we should mention that there are many instances of lost opportunities of synergy between general nonlinear potential theories and nonlinear PDEs, with many concepts being discovered independently in the two realms. We will focus on one such example in the discussion of the Jensen and Słodkowski Lemmas towards the end of this introduction. Another major aim of this work is to contribute to the many opportunities for productive interplay between these two realms. The reader may also want to consult the introductions of [8] and [32] for a more organic discussion.
@@
Before describing the contents of the present work, it will be useful to recall the definitions of subequation (constraint sets) and their subharmonics which are the basic objects in any potential theory. A subequation on an open subset $X\subset{\mathbb{R}}^{n}$ is a (closed) subset
$\mathcal{F}\subset\mathcal{J}^{2}(X):=X\times\mathcal{J}^{2}:=X\times{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$
@@
of the bundle of $2$-jets over $X$ satisfying the properties of positivity (P), negativity (N) and topological stability (T). Properties (P) and (N) are monotonicity properties which combine to ask that for each $x\in X$ the fibers $\mathcal{F}_{x}:=\{J\in\mathcal{J}^{2}:(x,J)\in\mathcal{F}\}$ satisfy
@@
these monotonicity properties correspond to the differential operator being proper elliptic. This is the minimal monotonicity that must be assumed for the validity of the natural comparison principles associated to both $\mathcal{F}$ and $F$. Property (T) helps to ensure a robust potential theory associated to a subequation $\mathcal{F}$ (see Definitions 5.9 and 5.5) including a natural notion of duality (see Section 6). Given a subequation $\mathcal{F}$, a function $u\in C^{2}(X)$ is said to be $\mathcal{F}$-subharmonic in $x\in X$ if
$J^{2}_{x}u:=(u(x),Du(x),D^{2}u(x))\in\mathcal{F}_{x}$ (0.1)
and $\mathcal{F}$-subharmonic on $X$ if (0.1) holds for each $x\in X$. For merely upper semicontinuous functions $u\in\operatorname{USC}(X)$, the definition (0.1) is taken in the viscosity sense; that is,
$J^{2,+}_{x}u\subset\mathcal{F}_{x}$ (0.2)
@@
$J^{2,+}_{x}u:=\{J^{2}_{x}\varphi:\varphi\ \text{is $C^{2}$ near $x$ and}\ u\leqslant\varphi\ \text{near $x$ with equality at $x$}\}.$
Property (P) is necessary in order for (0.2) to be useful and should be regarded as the weakest form of ellipticity necessary for a meaningful viscosity theory of subharmonics and subsolutions, as explained in Remark 5.2. Property (P) also ensures coherence in the sense that for $u$ twice differentiable in $x$, the classical notion (0.1) implies the viscosity notion (0.2). Coherence is discussed in Remark 5.3. Duality is then used to define the notions of a function $u$ being $\mathcal{F}$-superharmonic and $\mathcal{F}$-harmonic (see Sections LABEL:sec:conset and 6).
We are now ready to give a description of the contents of this work. We will attempt to highlight the main results and to indicate how they relate to the preprints [25] and [26] on which this primer is based. The work is divided into two parts. Each part begins with a detailed guide for the reader, which complements the following synopsis of the two parts.
Part I is dedicated to three fundamental and complementary aspects of quasi-convex analysis. The first aspect concerns differentiability properties of first and second order for locally quasi-convex functions, including the essential Alexandrov’s Theorem 1.25 on the almost eveywhere differentiability of second order for (locally quasi-) convex functions. The second aspect concerns a detailed analysis of their upper contact jets and upper contact points and culminates in two fundamental results on locally quasi-convex functions at or near points of upper (quadratic) contact: the Upper Contact Jet Theorem 2.12, which describes first and second order differentiability properties and the Summand Theorem 2.16 which deduces useful and important information about upper contact points and upper contact jets of two functions $u,v$ from an upper contact jet for the sum $w(x)=u(x)+v(x)$. This last result has a clear relation to the Theorem on Sums of Crandall, Ishii, and Lions [13], which instead treats sums such as $w(x,y)=u(x)+v(y)-\alpha|x-y|^{\alpha}$ with $\alpha>0$ and large, and is a lynchpin on which much of the conventional viscosity theory rests. Interestingly, the Summand Theorem yields immedately a quasi-convex version of the Theorem on Sums (as will be noted in Corollary 2.19). The third aspect concerns the deep analytical results which ensure the existence of sets of contact points with positive measure. This part of the analysis will culminate in the so-called Jensen–Słodkowski Theorem 3.28, which unifies and generalizes two historically distinct approaches to a central problem in a viscosity approach in potential theory and operator theory. We will return to this point at the end of the introduction as an illustration of parallel developments in nonlinear potential theories and nonlinear PDEs.
These three aspects of Part I are essential for the implementation of the viscosity theory of subharmonics and supersolutions of Part II, as will be described at the end of the guide for the reader to Part I. Moreover, they represent the “hard analysis” part of the viscosity theory. This part expands upon results contained primarily in [25], but also expands upon some results in [26].
Part II is dedicated to general nonlinear second order potential theories and the fundamental role that quasi-convex functions play in such theories. The presentation will involve three stages which can be summarized in the following way. Stage one gives a self-contained introduction to general second order potential theories, which is the context in which we will operate. Particular emphasis will be paid to the notions of duality and monotonicity which will have important consequences in all that follows.
Stage two is represented by two additional fundamental results of [26]. The first result is the Almost Everywhere Theorem, which will also be called the AE Theorem, states that a locally quasi-convex function which is $\mathcal{F}$-subharmonic almost everywhere (with respect to Lebesgue measure) in the classical sense (0.1) is in fact $\mathcal{F}$-subharmonic everywhere in the viscosity sense (0.2). This is particularly relevant in light of the easy extension of Alexandrov’s theorem from convex to locally quasiconvex functions. The second result, which follows from the AE Theorem, is a quasi-convex version of the so-called Subharmonic Addition Theorem 8.3. Subharmonic addition theorems concern the validity ot the following implication: given a trio $\mathcal{F},\mathcal{G}$ and $\mathcal{H}$ of subequation constraint sets in $\mathcal{J}^{2}(X)$ on an open set $X\subset{\mathbb{R}}^{n}$ (which then determine spaces of subharmonic functions $\mathcal{F}(X),\mathcal{G}(X)$ and $\mathcal{H}(X)$ on $X$), from a purely algebraic statement on the space of $2$-jets
$\forall\,x\in X:\quad\mathcal{F}_{x}+\mathcal{G}_{x}\subset\mathcal{H}_{x}\quad\text{({\em jet addition})},$ (0.3)
one concludes an analytic statement at the potential theory level
$\mathcal{F}(X)+\mathcal{G}(X)\subset\mathcal{H}(X)\quad\text{({\em subharmonicaddition})}.$ (0.4)
@@
When the subequations have constant coefficients, the Subharmonic Addition Theorem is extended in Corollary 8.9 to upper semicontinuous subharmonics by quasi-convex approximation, which preserves subharmonicity in the constant coefficient setting (see Proposition 8.8). Although we will not discuss this here, in the variable coefficient case, the aforementioned assumption of fiberegularity of the subequations yields a suitable alternative quasi-convex approximation scheme and the Subharmonic Addition Theorem extends to this setting (see [11] for details).
@@
Stage three concerns the use of quasi-convex analysis in the proof of comparison principles for nonlinear potential theories, as sketched in [26]. Two distinct settings will be discussed. The first setting is when there is sufficient monotonicity in the sense that the subequation $\mathcal{F}$ defining the potential theory admits a constant coefficient monotonicity cone subequation $\mathcal{M}$; that is, (0.3) holds with $\mathcal{G}=\mathcal{M}$ and $\mathcal{H}=\mathcal{F}$ and duality gives the jet addition formula
$\forall\,x\in X:\quad\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\mathcal{M}.$

### v3

# Kevin R. Payne Dipartimento di Matematica “F. Enriques” Università degli Studi di Milano Via C. Saldini 50

@@
[name=subject,title=Subject Index] [name=symbol,title=Symbol Index]
A primer on semiconvex functions in general potential theoriesfootnotetext: Revised version of the work A primer on quasi-convex functions in nonlinear potential theories, with modified title, reorganization of introductory material and rewritten last chapter. Preprint version of content to be published in Springer Lecture Notes in Mathematics, vol. 2371.
Kevin R. Payne Dipartimento di Matematica “F. Enriques”
Università degli Studi di Milano
Via C. Saldini 50
20133 Milano, Italy
kevin.payne@unimi.it (Kevin R. Payne)
 and  Davide Francesco Redaelli Dipartimento di Matematica
Università degli Studi di Roma “Tor Vergata”
Via della Ricerca Scientifica 1, 00133 Roma, Italy
redaelli@mat.uniroma2.it (Davide Francesco Redaelli)
(Date: April 15, 2025)
Key words and phrases:
semiconvexity, upper contact jets, subequation constraint sets, generalized subharmonics, viscosity solutions, comparison principles, fully nonlinear degenerate elliptic operators
2010 Mathematics Subject Classification:
26B25, 26B05, 31-01, 31C35, 35B51, 35D40
Preface
This book gives a detailed and self-contained treatment of the fundamental role that semiconvex functions play in general (nonlinear second order) potential theories. Such potential theories are concerned with the study of generalized subharmonics; that is, the class of functions whose partial derivatives up to second order are constrained to belong to a suitable closed subset (called a subequation) of the space of $2$-jets. The boundary of a subequation provides a second order partial differential equation. In any potential theory, there is a natural Dirichlet problem for generalized harmonics with prescribed boundary values. There are two natural notions of subharmonics in any potential theory: the classical notion for twice differentiable functions and the viscosity notion for upper semicontinuous functions, in which the classical second order differential at each point is replaced by the set of all upper contact jets of second order.
The utility of the viscosity theory stems from two considerations. First, a harmonic function in a general (nonlinear) potential theory may not be twice differentiable in the classical sense. Second, the viscosity notion has many stability properties which are extremely useful in limiting constructions, such as those needed for existence results by way of Perron’s method and for uniqueness results by way of the comparison principle (which is also needed for Perron’s method).
Semiconvex functions are used to build a bridge between the classical and viscosity notions of solutions to the natural Dirichlet problem in any potential theory. Since semiconvex functions are (by definition) quadratic perturbations of convex functions, they inherit the differentiability properties of convex functions, such as second order differentiability (in the Peano sense) on sets of full Lebesgue measure (Alexandrov’s theorem). More is true. There are deep and underused properties of their upper contact jets (essential for the viscosity theory) which makes the class of semiconvex functions a useful waystation when investigating potential theoretic results. In particular, the class of functions which are $\lambda$-semiconvex for some $\lambda\geqslant 0$ are the subharmonics of a certain potential theory. Moreover, as is well known, one can (locally) approximate upper semicontinuous functions which are (locally) bounded above with semicontinuous functions by way of the sup-convolution. This often allows one to transfer the wealth of information on upper contact jets for the semiconvex approximations to the upper semicontinuous limits. This deep and “hard” part of the viscosity theory is a central focus of this book. A brief discussion of the highlights of this analysis will be discussed in the introduction.
While general potential theories are interesting in and of themselves, the results presented here take on greater significance in light of a rapidly developing and fruitful interplay between general potential theories and the operator theory of fully nonlinear elliptic partial differential equations. Historically, research on the complex Monge-Ampère equation has had a strong connection to complex pluripotential theory. Beginning with a trio of papers in 2009 by Harvey and Lawson [24, 25, 26], this interplay is being applied with much success in order to embed a given partial differential operator into an associated potential theory and then to deduce properties about the operator theory from the corresponding potential theoretic properties. Moreover, these properties are shared by any operator which is compatible with the subequation that determines the potential theory. This often simplifies and unifies the operator theory. Results include, but are not limited to, the validity of the comparison principle (which implies uniqueness for the Dirichlet problem) and the application of Perron’s method for existence (which makes use of the comparison principle as well as a suitable notion of boundary pseudoconvexity that has an elegant potential theoretic formulation). The reader is invited to consult [9, 10, 11] and [12] as well as the organic survey presented in [35]. We will give a short history of this program at the end of the overview in the introduction. In that light, this book can be viewed as a prequel to the above mentioned works and our main aim is to provide a comprehensive presentation of the foundations on which this theory rests.
An important remark on terminology is in order. With some reservation, we have decided to use the more conventional term semiconvex for a function $u$ (defined on a convex subset of Euclidian space) such that $u+\frac{\lambda}{2}|\cdot|^{2}$ is a convex function for some $\lambda\geqslant 0$. This is equivalent to asking that the Hessian of $u$ is bounded below (in the viscosity sense). In the works mentioned above, the less conventional term quasi-convex is used, which is consistent with the use of quasi-plurisubharmonic in several complex analysis, as noted in [35]. However, in convex analysis, quasi-convexity is also used to mean that inverse images of half-lines bounded above are convex sets. In an attempt to avoid confusion over a wide audience, we have adopted the more conventional term semiconvex, as used in much of the viscosity literature, such as [14]. We would like to stress however that semiconvex functions are not “half-convex” in the same way that semicontinuous functions are “half-continuous”. Moreover, any $C^{2}$ function is locally semiconvex, which shows that semiconvex functions are far from being “half-convex”. Perhaps a different term such as “convexable” might be advisable, or to use Hessian bounded below as in [48].
Some declarations of intent are also in order. The present work was born with the purpose of unifying and harmonizing the two articles [28] and [29] of Harvey and Lawson on semiconvex functions and their use in general potential theories, which were deposited in the arXiv in 2016. These two preprints are full of foundational aspects of non-smooth analysis, some well-known, others less well-known and many new things. The authors of this work are extremely grateful to Harvey and Lawson for their encouragement to attempt this synthesis, which was the basis of the Master’s Thesis of the second author, supervised by the first author.
This book is more than a mere merging of the two preprints [28] and [29], whose combined length is less than one third of the present work. Beyond the question of length, the current presentation benefits from the numerous developments in the theory over the past decade. A discussion of what is “new” here can be found in the commentary section of the introduction.
This book is aimed at a wide audience, including professional mathematicians working in fully nonlinear PDEs as well as masters and doctoral students with interest in mathematical analysis. On the one hand, the book collects a wealth of results (some classical and some new) on how to treat second order differentiability in a pointwise manner for functions which are merely semicontinuous and the (viscosity) subharmonics in general potential theories. As such, we believe that much of this material should be more widely known by working mathematicians. On the other hand, much of the book has been tested in the classroom in the context of a master’s level course on the foundations of the viscosity theory for fully nonlinear elliptic PDEs.
We note that very little prior knowledge is necessary to read this book, which aims to be relatively self-contained. However, some familiarity with basic notions of mathematical analysis, such as the standard differential calculus of several variables, Lebesgue measure and integral, the notions of limsup and liminf are assumed. Good background references are [21] and [52], and also [19] for many advanced topics. In addition, some familiarity with harmonic functions and maximum principles for elliptic equations is certainly useful, such as what one finds in [18]. For further complementary reading, other than the papers mentioned in the book, we mention the classic monographs on convexity [38] and [47], on general potential theories [9], and elliptic equations [22], [7] and [41].
The authors express their deep gratitude to Blaine Lawson and Reese Harvey for their generosity and encouragement to perform and updated and enriched synthesis of the unpublished manuscripts [28] and [29]. In addition, numerous useful conversations with Reese Harvey have contributed to the development of this work. The authors are partially supported by the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM), Italy. Redaelli is partially supported by the Fondazione Cassa di Risparmio di Padova e Rovigo, Italy and by the Italian (EU Next Generation) PRIN project 2022W58BJ5.
Milano, Roma Kevin R. Payne
March 2025 Davide Francesco Redaelli
@@ Introduction
As described in the preface, this book is dedicated to foundational aspects of general (nonlinear second order) potential theories and fully nonlinear elliptic PDEs. In particular, we will systematically develop the fundamental role played by semiconvex functions as a bridge between the classical and viscosity theory of generalized subharmonics determined by a given subequation (constraint set) in the bundle of $2$-jets over open subsets of Euclidian spaces. In this introduction, we will first give an overview of the book, describing the organization of the presentation, the main concepts and important results that are contained within. Then we will give a commentary, which aims to place the book in the context of the rapidly developing and fruitful interplay between the general potential theories and the operator theory of elliptic PDEs. In addition, remarks concerning what is “new” in this book will be presented. The reader is invited to read the preface before beginning to look at the rest of the book and may want to use this introduction as guide to consult before or after having delved into the book.
Overview
The book is divided into two parts and also contains two appendices which complement the first part. In this section, we will present a synopsis of the contents.
Part I, which consists of Chapters 1—4, is dedicated to four fundamental and complementary aspects of semiconvex analysis. The first aspect (treated in Chapter 1 concerns differentiability properties of first and second order for locally semiconvex functions, including the essential Alexandrov’s Theorem 1.29 on the almost everywhere differentiability of second order for (locally semi-) convex functions. A complete and self-contained proof of Alexandrov’s theorem is provided in Appendix A and depends on two key ingredients: several important properties of the Legendre transform and a Lipschitz version of Sard’s theorem. For completeness, the highly technical proof of this last ingredient is also given in Appendix B.
The second aspect (treated in Chapter 2) concerns a detailed analysis of upper contact jets and upper contact points of locally semiconvex functions. This investigation culminates in two fundamental results on locally semiconvex functions at or near points of upper (quadratic) contact: the Upper Contact Jet Theorem 2.12 and the Summand Theorem 2.16. The first result groups together three fundamental properties of first and second order differentiability of a locally semiconvex function at or near a point of upper (quadratic) contact. The second result produces interesting and important information about upper contact points of a pair of locally semiconvex $u$ and $v$ at or near a point of upper (quadratic) contact of the sum $w=u+v$. Interestingly, the Summand Theorem immediately yields a semiconvex version of the Theorem on Sums (as will be noted in Corollary 2.19). The celebrated semicontinuous version of the Theorem on Sums is obtained after doubling variables; that is, by considering $w(x,y)\vcentcolon=u(x)+v(y)$ and is the linchpin on which much of the conventional viscosity theory of Crandall-Ishii-Lions [14] rests. For completeness, in Section 4.3, we will also give a novel proof of the semicontinuous version as a corollary of the semiconvex version and the standard device of semiconvex approximation by way of the sup-convolution which produces a controllable displacement of the upper contact points of every given upper contact jet. Elementary properties of this approximation are discussed in Section 4.2.
The third aspect (treated in Chapter 3) concerns additional deep analytical results needed for a robust viscosity theory (which is based on quadratic upper test functions). One has the central problem of how to ensure a priori the existence of a sufficient amount of upper contact points for an upper semicontinuous function. This is accomplished by combining Alexandrov’s Theorem, the sup-convolution approximation of semicontinuous functions by semiconvex functions and the results presented in this chapter. The results of this Chapter ensure the existence of sets of contact points with positive Lebesgue measure nearby points of strict upper contact of locally semiconvex functions. This part of the analysis will culminate in the so-called Jensen–Słodkowski Theorem 3.27, which unifies and generalizes two historically distinct approaches to the central problem in a viscosity approach in potential theory and operator theory. The most well-known result of this type is due to Jensen [39] and is recalled in Jensen’s lemma 3.6. This result says that nearby a strict local maximum point of a semiconvex function $w$ there is a set of positive measure for which small linear perturbations of $w$ still has a local maximum. An older result of Słodkowski [48], is recalled in Słodkowski’s density estimate 3.2 and concerns a lower bound on the Lebesgue density of sublevel sets of the largest eigenvalue of the Hessian of a convex function. An important and impressive insight of Harvey and Lawson [28] is that if one translates these two results into the language of upper contact points, these results are equivalent. We state these reformulations as the HL–Jensen Lemma 3.22 and the HL–Słodkowski Lemma 3.21 and prove that they are equivalent in Proposition 3.23. Moreover, these two reformulations can be merged into the aforementioned Jensen–Słodkowski Theorem. This story represents one of the most interesting and unique aspects of this book. We will return to this point at the commentary section below as an illustration of parallel developments in general potential theories and nonlinear elliptic PDEs.
Chapter 3 also includes various consequences of the Jensen–Słodkowski Theorem including the proof of the partial upper semicontinuity of second derivatives which is the most delicate part of the aforementioned Upper Contact Jet Theorem as well as a semiconvex version of Słodkowski’s Largest Eigenvalue Theorem 3.29. Moreover, in the last section of the chapter, an alternative proof of the Jensen–Słodkowski Theorem is given which only makes use of an Area Theorem 3.34 for gradients of locally semiconvex functions and the resulting Alexandrov’s maximum principle 3.35 for locally semiconvex functions.
The fourth and final apect is treated in Chapter 4 and concerns the essential role that semiconvex functions play in the viscosity theory of semicontinuous functions by way of the aforementioned semiconvex approximation of upper semicontinuous functions (see Theorem 4.10). One also finds an alternative proof of the conventional Theorem on Sums of Crandall–Ishii–Lions which makes use of the upper contact jet technology developed up to this point.
These four aspects of Part I are essential for the implementation of the viscosity theory of subharmonics and supersolutions of Part II, which will be described next. We repeat the important assertion that the ingredients of Part I represent the “hard analysis” part of the viscosity theory.
@@
To facilitate the synopsis of Part II, it will be useful to anticipate the definitions of a subequation (constraint set) and their subharmonics which are the basic objects in any potential theory. A subequation on an open subset $X\subset{\mathbb{R}}^{n}$ is a (closed) subset
$\mathcal{F}\subset\mathcal{J}^{2}(X)\vcentcolon=X\times\mathcal{J}^{2}\vcentcolon=X\times{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$
@@
of the bundle of $2$-jets over $X$ satisfying the properties of positivity (P), negativity (N) and topological stability (T). Properties (P) and (N) are monotonicity properties which combine to ask that for each $x\in X$ the fibers $\mathcal{F}_{x}\vcentcolon=\{J\in\mathcal{J}^{2}:(x,J)\in\mathcal{F}\}$ satisfy
@@
[…]
$J^{2}_{x}u\vcentcolon=(u(x),Du(x),D^{2}u(x))\in\mathcal{F}_{x}$ (0.1)
[…]
$J^{2,+}_{x}u\subset\mathcal{F}_{x}$
@@
$J^{2,+}_{x}u\vcentcolon=\{J^{2}_{x}\varphi:\varphi\ \text{is $C^{2}$ near $x$ and}\ u\leqslant\varphi\ \text{near $x$ with equality at $x$}\}.$
[…]
$\mathcal{F}_{x}+\mathcal{G}_{x}\subset\mathcal{H}_{x},\ \ \forall\,x\in X\implies\mathcal{F}(X)+\mathcal{G}(X)\subset\mathcal{H}(X).$ (0.2)
@@
When the subequations have constant coefficients, the Subharmonic Addition Theorem is extended in Corollary 8.9 to upper semicontinuous subharmonics by semiconvex approximation, which preserves subharmonicity in the constant coefficient setting (see Proposition 8.8).
Although we will not discuss this in the book, in the variable coefficient setting an additional assumption of fiberegularity of the subequations yields a suitable alternative semiconvex approximation scheme and the Subharmonic Addition Theorem extends to the variable coefficient setting (see [12] for details). The notion of fiberegularity is explained in the Commentary towards the end of this introduction.
@@
Stage three is presented in Chapter 9 and concerns the use of semiconvex analysis in the proof of comparison principles for nonlinear potential theories. Two distinct settings will be discussed. The first setting (treated in Section 9.2) concerns the case of sufficient monotonicity in the sense that the subequation $\mathcal{F}$ defining the potential theory admits a constant coefficient monotonicity cone subequation $\mathcal{M}$. This means that there exists a subequation $\mathcal{M}$, with constant fibers that are a closed convex cone in $\mathcal{J}^{2}$ (with vertex at the origin), such that
[…]
that is, the jet addition formula of (0.2) holds with $\mathcal{G}=\mathcal{M}$ and $\mathcal{H}=\mathcal{F}$. Duality then gives the jet addition formula
$\forall\,x\in X:\quad\mathcal{F}_{x}+\widetilde{\mathcal{F}}_{x}\subset\widetilde{\mathcal{M}}.$